# Partial persistence of memory in bubble breakup: incomplete universality acquired by broken symmetry

### v8

@@ II.1 The incomplete university: partial persistence of memory
The independence of the master curve representing the self-similarity from the length scales $D_{0}$ and $R$ shown in Fig. 5 (a) and dependence on $e$ shown in Fig. 6 (a3) reveal a new category of the incomplete universality for the memory of singular dynamics discussed in Introduction. This is because these lengths are the scales that set the initial boundary conditions of the present problem. In the present case, the master curve loses the memory on the scales $D_{0}$ and $R$, but retains the memory of the smallest scale $e$, which is quite natural from the viewpoint of scale separation, with remaining dependence on the material parameter on $\Delta\rho$ (and independence from $\eta$).
II.2 The pre- and post-breakup dynamics
In the case of the viscous-capillary breakup of a bubble under no geometrical constraint, where the axisymmetric liquid-air interface is described by $r=h(z,t)$ in the cylindrical coordinate $(r,\theta,z)$, the non-universal self-similar dynamics could directly continue from the pre- to post-breakup dynamics (see, e.g., Sec. 3.5.1 of [5]): the same self-similar structure is maintained, meaning that the factors for rescaling to achieve collapse onto the master curve are the same when expressed as a function of a dimensionless time $|\widetilde{t}^{\prime}|\sim\left|t-t_{c}\right|$. In other words, the pre- and post-breakup are both expressed as $h(z,t)=|\widetilde{t}^{\prime}|\Gamma(z/|\widetilde{t}^{\prime}|^{2})$.
As for the confined and non-axisymmetric breakup of the present focus, the pre-breakup dynamics was studied in the previous studies [14, 15] and two distinct cases were reported. When $D_{0}>\kappa^{-1}=\sqrt{\gamma/\rho g}\simeq 1.8$ mm, as in the present case, the liquid-sheet is formed at the constriction point, which leads to the breakup. When $D_{0}$ is smaller, the axisymmetric corn is formed which detaches from the disk without breakup. In both cases, the dynamics was found to be self-similar and the self-similar structure of the pre-breakup corresponding to the present case of $D_{0}>\kappa^{-1}$ was reported as $h(z,t)=2|\widetilde{t}^{\prime}|\Gamma(z/|\widetilde{t}^{\prime}|)$, which is significantly different from Eq. (2), i.e., $h(z,t)\simeq|\widetilde{t}^{\prime}|^{2\Delta^{\prime}}\widehat{\Gamma}(z/|\widetilde{t}^{\prime}|^{\Delta^{\prime}})$. This suggests the scenario for the continuation from the pre- to post-breakup could be quite different. In addition, for the pre-breakup, although the master curve for $D_{0}>\kappa^{-1}$ is shown to be independent of $D$, $R$, and $\eta$ in a certain range [14], the dependence on $e$ has yet to be confirmed. The problem of the continuation and the dependence on $e$ for the pre-breakup for $D_{0}>\kappa^{-1}$ will be discussed elsewhere in the near future.
II.3 Physics at the level of dimensional analysis
We may regard the present problem as finding a solution for Navier-Stokes equation for a viscous liquid, neglecting the role of air. The boundary conditions may involve, in addition to $R$, $e$, and $D$, the velocity of the tip at $t=t_{c}$, if we start to solve the post-breakup dynamics from $t=t_{c}$. The tip velocity is known to scale with the falling velocity of the disk in the previous study on the pre-breakup dynamics [14]: The velocity is shown to scale as $v_{G}\sim\Delta\rho gD^{2}/\eta$, as a result of the balance of gravitational and viscous energy (per time) for the disk: $\Delta\rho gR^{2}Dv_{G}\sim\eta(v_{G}/D)^{2}R^{2}D$ (for $D\simeq D_{0}$). In addition, we see in Eq. (2) and (1) any dependence on $\gamma$ but on $\eta$, $g$, and $\Delta\rho$. Thus, we expect
$h=f(t^{\prime},z,\Delta\rho,\eta,g,R,e,D)$ (5)
we have 9 dimensional variables, of which only 6 are independent, since the dimension of the unit of all variables can be derived from the three fundamental units, kg, m, and s.
@@ II.3 Physics at the level of dimensional analysis
From the Buckingham $\pi$ theorem (see, e.g., Appendix C of [5]), we expect a relation $\pi_{0}=\Xi(\pi_{1},\pi_{2},\ldots,\pi_{5})$, where $\pi_{i}$’s are 6 independent dimensional variables and $\Xi$ is a dimensionless function. We select these dimensionless variables as follows. A natural characteristic scale in the $z$ direction is $z_{m}$, from which we define $\pi_{1}=z/z_{m}-1$. A natural unit $h^{\ast}$ (in the $x$ direction) for $h$ can be introduced through a curvature relation, $1/R\sim h^{\ast}/z_{m}^{2}$, from which we set $\pi_{0}=h/h^{\ast}$. To select the remaining 4 independent variables, we focus on 4 length scales: $l=\eta/(\Delta\rho gt^{\prime})$, $R$, $e$, and $D$, which are normalized by $z_{m}$ to determine 4 dimensionless variables, $\pi_{2},\ldots,\pi_{5}$. In this way, to be consistent with our experiment, we may arrive at
$h=\frac{z_{m}^{2}}{R}\Xi((z-z_{m})/z_{m},l/z_{m},R/z_{m},D/z_{m},e/z_{m})$ (6)
Here, we may expect that near the breakup point where $z_{m}\sim(t^{\prime})^{\Delta^{\prime}}$ with $10<\Delta^{\prime}<$ 1 is small so that the right-hand side of the equation becomes independent of the second to the fourth dimensionless variables (with $l/z_{m}\rightarrow\infty$ and $R/z_{m},D/z_{m}\rightarrow 0$ but with a finite $e/z_{m}$), by which Eq. (2) with remaining dependence on $e$ is reproduced.
II.4 Importance of exploring symmetry
@@
In critical phenomena, where universality plays a significant role as in the singular dynamics [16, 17] and such universality emerges because the physics at large scales becomes so important, a myriad of universality classes have been found by exploring symmetry and dimensionality, which has propelled the developments of modern physics in different sub-domains such as soft and hard condensed matter, non-equilibrium systems, and active matter [18, 19, 20, 21]. Given this, the present study suggests a promising direction for the study on the singular dynamics widely observed in nature by exploring the symmetry in confined geometries and by drawing an analogy with critical phenomena. The present findings on the singular dynamics with broken axisymmetry have implications for understanding many natural phenomena and industrial processes involving confinement, encountered in a various fields, from geology and petroleum industry to microfluidics with applications such as in medicine and biochemistry [22, 23, 24, 25].
Acknowledgements.
@@
This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596.
Appendix A Experimental Details
The setup and experimental procedure is the same as those in the previous study [14] not on the post-breakup dynamics of the present focus but on the pre-breakup. We here summarize important points to realize high reproducibility. We need to achieve (1) the thickness $e$ of the lubrication films on both surfaces of the disk to be equal, i.e., $e=(D-D_{0})/2$, as indicated in Fig. 1 (a), (2) the initial velocity of the metal disk to be zero at the entry, and (3) the contact angle of the surface of the disk to be zero. Accordingly, (1) we set a gate with the gap $D_{0}$ at the top of cell by placing two small plates of thickness $e$ on the inside surfaces of the cell, (2) we fall the disk so that the bottom of the disk is initially in contact with the air-liquid interface, and (3) we wipe the oil on the surface of the disk after dipping to coat the surface by a thin layer of the oil. To capture the dynamics, we analyzed the images with Image J and self-made Python codes, which were obtained with a high-speed camera (FASTCAM Mini UX 100, Photron) with a lens (Micro NIKKOR 60 mm f2.8G ED, Nikon) at 1000 to 2000 frames per second (fps).
Appendix B Definitions of the shape function $h$ and setting of space-time coordinates
@@
The inner edge by which we define the shape corresponds to the contact line on the front surface of the back cell wall, whose surface is totally wetting (thin layer of oil exists ahead of ”the contact line”). The relation $x=h(z,t)$ represents the shape outlined by the contact line on the back cell plate.
@@
The critical time $t_{c}$ used in the analysis was determined as follows. Towards breakup, the constriction region starts to form a thin thread (when seen from the front), which finally pinches off, by which moment, we define $t=t_{c}$. Precisely speaking, the pinch-off is judged not by the inner edge but the outer edge as indicated in Fig. 1 (b) left (the inner edge breaks up before the outer edge), but the positions $z_{c}$ and $z_{m}$ are determined by the inner edge (In Fig.1 (b), these positions are actually slightly above the dashed horizontal lines by the amount of the thickness of the line representing the interface). In addition, we set the time label 00 ms as the snapshot just before the pinch off, using snapshots obtained at 1000 frames per second, which means there could be a difference at most 1 ms between our time label 0 ms and the actual critical time $t=t_{c}$. Because of this possible error in determining the origin of time, we avoid using snapshots too close to $t=t_{c}$ and use only those at $t\gtrsim 10$ ms.

### v9

@@ Abstract
The incomplete universality: partial persistence of memory— The independence of the master curve from the length scales $D_{0}$ and $R$ shown in Fig. 5 (a) and dependence on $e$ shown in Fig. 6 (a3) reveal a new category of the incomplete universality for the memory of singular dynamics discussed in Introduction. This is because these lengths are the scales that set the initial boundary conditions of the present problem. In the present case, the master curve loses the memory on the scales $D_{0}$ and $R$, but retains the memory of the smallest scale $e$, which is quite natural from the viewpoint of scale separation, with remaining dependence on the material parameter on $\Delta\rho$ (and independence from $\eta$).
As a matter of fact, Fig. 6 (a3) shows that the two curves obtained for different $e$’s but the same $\Delta\rho$ become indistinguishable near the tip (although they seem distinguishable from the other obtained for a different $\Delta\rho$). This observation is in accord with our result of the two-step fitting: $(a,b)=(0.9\pm 0.1,0.6\pm 0.05)$ for Fig. 6 (b1) and $(1.9\pm 0.2,0.7\pm 0.05)$ for Fig. 6 (b2): $a=1.9$ is indistinguishable from $a=1.8$ (although $a=0.9$ is distinguishable). This observation that near the tip universality tends to recover is also expected from separation of scales: very close to the tip, the characteristic length for the collapse tends to become small and thus become well separated even from the smallest length scale $e$.
[…]
Physics at the level of dimensional analysis— We may regard the present problem as finding a solution for Navier-Stokes equation for a viscous liquid, neglecting the role of air, giving initial BCs. It would be easier to give the initial BCs at the time $t=t_{c}$ rather than the time of the entry of the disk because after $t=t_{c}$ we can forget about the disk motion by giving the velocity of the tip of air at $t=t_{c}$. This velocity is given as $v_{G}\sim\Delta\rho gD^{2}/\eta$ by balancing gravitational and viscous energy (per time) for the disk, $\Delta\rho gR^{2}Dv_{G}\sim\eta(v_{G}/D)^{2}R^{2}D$ (for $D\simeq D_{0}$), in the previous study on the pre-breakup dynamics [15]. In addition, from Eqs. (2) and (1) with the dependence on $e$ of the master curve, we expect the BCs involves $\Delta\rho,\eta,g,R,e,$and $D$ but not on $\gamma$. Thus, we expect $h=f(t^{\prime},z,\Delta\rho,\eta,g,R,e,D)$. Here, we have 9 dimensional variables, of which only 6 are independent, since the dimension of the unit of all variables can be derived from the three fundamental units, kg, m, and s.
@@
From the Buckingham $\pi$ theorem (see, e.g., Appendix C of [6]), we expect a relation $\pi_{0}=\Xi(\pi_{1},\pi_{2},\ldots,\pi_{5})$, where $\pi_{i}$’s are 6 independent dimensional variables and $\Xi$ is a dimensionless function. We select these dimensionless variables as follows. A natural characteristic scale in the $z$ direction is $z_{m}$, from which we define $\pi_{1}=z/z_{m}-1$. A natural unit $h^{\ast}$ (in the $x$ direction) for $h$ can be introduced through a curvature relation, $1/R\sim h^{\ast}/z_{m}^{2}$, from which we set $\pi_{0}=h/h^{\ast}$. To select the remaining 4 independent variables, we focus on 4 length scales: $l=\eta/(\Delta\rho gt^{\prime})$, $R$, $e$, and $D$, which are normalized by $z_{m}$ to determine 4 dimensionless variables, $\pi_{2},\ldots,\pi_{5}$. In this way, to be consistent with our experiment, we may arrive at $h=\frac{z_{m}^{2}}{R}\Xi((z-z_{m})/z_{m},l/z_{m},R/z_{m},D/z_{m},e/z_{m})$. Here, we may expect that near the breakup point where $z_{m}\sim(t^{\prime})^{\Delta^{\prime}}$ with $0<\Delta^{\prime}<1$ is small so that the right-hand side of the equation becomes independent of the second to the fourth dimensionless variables (with $l/z_{m}\rightarrow\infty$ and $R/z_{m},D/z_{m}\rightarrow 0$ but with a finite $e/z_{m}$).
In this way, the dimensional analysis provides a natural understanding of the scaling structure in Eq. (2) with the function $\Gamma$ dependent on $e$ but not on $R$ and $D$, based on which structure we proposed the novel scenario of the incomplete universality. Although the present analysis cannot determine the scaling exponents (such cases are known as the self-similarity of the second kind [5]), a renormalization group analysis recently developed for the non-confined bubble breakup [16] will be promising to determine the exponents and to elucidate the origin of the scaling crossover in the present study.
Importance of exploring symmetry— We stress here that it is becoming clear that the present experimental system that can break the axisymmetry contains extremely rich physics. In the previous studies [15, 17, 18], three distinct regimes have been found by changing the range of confining length scales: the sheet-forming regime with breakup, the corn-forming regime with and without breakup. In the first and second, a sheet and a corn of air (which are both non-axisymmetric) are respectively formed at the constriction point leading to breakup, while, in the third, a corn detaches from the disk without appearance of the constriction point. In the third, in particular, analogy with critical phenomena is deeply explored to find the exponents are dependent on a length scale, which is a novel feature not found in the first and second regimes. In all of the three studies [15, 17, 18], we focus on the pre-detachment dynamics where detachment includes breakup. In the present study, we focus on the post-breakup for the first time, while, experimentally, the post-breakup dynamics has not been explored even in the non-confined case. As a result, we found a self-similar structure significantly different from those in pre-breakup dynamics as seen above and successfully obtained the data strongly supports the novel scenario of the incomplete universality.
@@
The remarkable physical richness of the present experimental system was uncovered by breaking of symmetry, which reminds the importance of symmetry and dimensionality in critical phenomena [19, 20]: a myriad of universality classes have been found by exploring symmetry and dimensionality, which has propelled the developments of modern physics ranging from soft and hard condensed matter, non-equilibrium systems to active matter [21, 22, 23, 24]. Given this, the present study suggests a promising direction for the study on the singular dynamics: exploring the symmetry in confined geometries with drawing an analogy with critical phenomena. This direction is all the more promising if we remind that the singular dynamics has been widely observed and confinement is involved in many cases of natural phenomena and industrial processes ranging from geology and petroleum industry to microfluidics with applications such as in medicine and biochemistry [25, 26, 27, 28].
@@
Acknowledgments— This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596.
Appendix
@@
A1: Definitions of the shape function $h$ The inner edge by which we define the shape corresponds to the contact line on the front surface of the back cell wall, whose surface is totally wetting (thin layer of oil exists ahead of ”the contact line”). The relation $x=h(z,t)$ represents the shape outlined by the contact line on the back cell plate.
@@
A2: Setting of space-time coordinates— The critical time $t_{c}$ used in the analysis was determined as follows. Towards breakup, the constriction region starts to form a thin thread (when seen from the front), which finally pinches off, by which moment, we define $t=t_{c}$. Precisely speaking, the pinch-off is judged not by the inner edge but the outer edge as indicated in Fig. 1 (b) left (the inner edge breaks up before the outer edge), but the positions $z_{c}$ and $z_{m}$ are determined by the inner edge (In Fig.1 (b), these positions are actually slightly above the dashed horizontal lines by the amount of the thickness of the line representing the interface). In addition, we set the time label 0 ms as the snapshot just before the pinch off, using snapshots obtained at 1000 frames per second, which means there could be a difference at most 1 ms between our time label 0 ms and the actual critical time $t=t_{c}$. Because of this possible error in determining the origin of time, we avoid using snapshots too close to $t=t_{c}$ and use only those at $t\gtrsim 10$ ms.